Last month, Mrs. Rita bought x kilograms of watermelon for 400 pesos. Today she pay same amount but gets 2 kilograms less due to price increase of 10 pesos per kilogram. find the value of x and the original price of the watermelon
1
Expert's answer
2022-01-04T08:27:00-0500

Originally, there are "x" kilograms of watermelons. Let the price of watermelons per kilogram be "y". Then, "x\times y=400........(i)"

Currently, the amount of watermelons she buys is "(x-2)" kilograms. The price she pays per kilogram is "(y+10)" pesos. The total amount she pays is 400 pesos. Then, "(y+10)\times(x-2)=400.......(ii)"

From equation "(i),"

"x\times y=400\implies x={400\over y}"

Substituting for the value of "x" given above in equation "(ii)" we have,

"(y+10)\times({400\over y}-2)=400"

Opening the brackets we have,

"400-2y+{4000\over y}-20=400"

"{4000\over y}-2y=20"

Multiplying through by "y" and dividing through by 2,

"2000-y^2=10y\implies -y^2-10y+2000=0.......(iii)"

Equation "(iii)" is a quadratic equation which can be solved by using quadratic formula given as,

"y={-b\pm\sqrt{b^2-4ac}\over2a}"

where, "a=-1, \space b=-10,\space c=2000"

Therefore,

"y={10\pm\sqrt{(-10)^2-4(-1\times 2000)}\over 2(-1)}={10\pm\sqrt{8100}\over-2}={10\pm90\over-2}"

Solving for the value of "y",

"y={(10+90)\over -2}={100\over-2}=-50 \space or \space {(10-90)\over-2}={-80\over -2}=40"

Since the value of "y" can not be negative, we have that "y=40"

From equation "(i)", the value of "x" is,

"x\times y=400\implies x={400\over y}={400\over 40}=10"

Therefore, "x=10" and the original price of watermelon is "y=40" pesos per kilogram.